Resolvent estimates for plane Couette flow

Pablo Braz e Silva Universidade Federal de Pernambuco, Brazil

We discuss the problem of deriving estimates for the resolvent of the linear operator associated with three-dimensional perturbations of plane Couette flow and determining its dependence on the Reynolds number.